OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points (3/2, -2) and (-3,7/3)?

OpenStudy (gorv):

\[slope=\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (gorv):

\[\frac{ -2-7/3 }{ 3/2-(-3) }\]

OpenStudy (gorv):

\[\frac{ -13/3 }{ 3/2+3 }\]

OpenStudy (gorv):

\[\frac{ -13/3 }{ 9/2 }=\frac{ -13*2 }{ 9*3 }=\frac{ -26 }{ 27 }\]

OpenStudy (gorv):

@JerryK